A projectile is launched with speed Vo at an angle "theata" above the horizontal. The launch point is at a height h above the ground.
(a) show that if air resistance is neglected the horizontal range, R that the projectile travels before striking the ground is R = Vo*cos(theata)\g(VoSin"theat a"+√Vo^2Sin^2" theata"+2gh).
(b) what will be the expression for R if the launch point is at the ground level so that h = 0
(c)what
1
Expert's answer
2019-10-21T09:17:01-0400

a) The maximum height the projectile will reach after it's launched from the height "h":


"H=h+\frac{(v_0\text{ sin}\theta)^2}{2g}=\frac{2gh+(v_0\text{ sin}\theta)^2}{2g}."

And this will take time


"t_1=\frac{v_0\text{ sin}\theta}{g}."


Then it will fall back to the ground during


"t_2=\sqrt{\frac{2H}{g}}=\frac{\sqrt{2gh+(v_0\text{ sin}\theta)^2}}{g}."

Meanwhile the projectile will move forward with constant velocity


"v_x=v_0\text{ cos} \theta,"

and of course it will cover the range


"R=v_x(t_1+t_2)=\\=\frac{v_0\text{ cos} \theta}{g}(v_0\text{ sin}\theta+\sqrt{2gh+(v_0\text{ sin}\theta)^2})."

b) Substitute "h=0" and get simple relation for the horizontal range:


"R(0)=\frac{v_0^2\text{ sin} (2\theta)}{g}."

c) That is the answer to this problem.